### v2

# PAL: Probing Audio Encoders via LLMs - Audio Information Transfer into LLMs

Abstract
Integration of audio perception into large language models (LLMs) is an emerging research area for enabling machine listening applications, yet efficient transfer of rich audio semantics from audio encoders to LLMs remains underexplored. The most widely used integration paradigm projects the audio encoder output tokens into the LLM input space (e.g., via an MLP or a Q-Former), then prepends or inserts them to the text tokens. We refer to this generic scheme as Prepend to the LLM’s input token space (PLITS) integration. We propose an efficient alternative, Lightweight Audio LLM Integration (LAL). LAL introduces audio representations solely via the attention mechanism within different layers of the LLM, bypassing its feedforward module. LAL encodes rich audio semantics at an appropriate level of abstraction for integration into different blocks of LLMs. Our design significantly reduces computational overhead compared to existing integration approaches. Observing with Whisper that the speech encoder benefits from PLITS integration, we propose an audio encoder aware approach for efficiently Probing Audio encoders via LLM (PAL), which employs PLITS integration for Whisper and LAL for general audio encoders. Under an identical training curriculum, LAL consistently maintains performance or outperforms existing integration approaches across multiple base LLMs and tasks. For general audio tasks, LAL improvement is up to 30% over a strong PLITS baseline while reducing memory usage by up to 64.1% and increasing throughput by up to 247.5%. Furthermore, for general audio-music-speech LLM, PAL, performs on par with a fully PLITS integration-based system but with substantially improved computational and memory efficiency. Project page:https://ta012.github.io/PAL/
1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Grattafiori et al., 2024; Jiang et al., 2024; Liu et al., 2024a) have emerged as the foundational technology for natural language interaction with machines, demonstrating remarkable conversational fluency. Despite this success, their perceptual capabilities remain limited primarily to text, restricting their ability to understand the physical world. This limitation has inspired significant research into multi-modal LLMs (MLLMs), which expand traditional LLMs by integrating additional sensory modalities such as vision (Vision LLMs) (Liu et al., 2023; Templeton et al., 2024; Wang et al., 2024), audio (Large Audio Language Models (LALMs) or simply audio-LLMs) (Deshmukh et al., 2023; Gong et al., 2024; Tang et al., 2024; Ghosh et al., 2024; 2025a), and other inputs (Brohan et al., 2023; Thawkar et al., 2023) to foster more natural, intuitive, and effective human-machine interfaces.
Recent advances in audio representation learning have produced powerful encoders trained with self-supervised objectives  Huang et al. (2022); Alex et al. (2025) and multimodal supervised objectives (CLAP Elizalde et al. (2023); Wu et al. (2023). We argue that the primary function of an audio LLM is to query the audio encoder’s representations via natural language and return the information the user wants. This requires an integration mechanism that reliably transfers key audio information, including event cues and temporal dynamics, from the encoder into the LLM embedding and concept space; once received, the LLM can apply its reasoning and generation to answer queries from concise facts to nuanced creative text. Our work focuses on  optimizing this information transfer.
An audio-LLM typically consists of (i) a large language model (LLM), (ii) one or more audio encoders, and (iii) an integration strategy that connects encoder outputs to the LLM. Regarding the selection of audio encoders, prior audio-LLM work often performs additional pretraining of a chosen encoder to extend its capabilities (Goel et al., 2025; Chu et al., 2024b; Gong et al., 2024). For example, the Whisper encoder (Radford et al., 2023), originally trained for speech-to-text transcription, is further adapted for general audio event understanding in (Goel et al., 2025); the AST encoder (Gong et al., 2022a; 2021) is retrained for language alignment in (Gong et al., 2024; Ghosh et al., 2024). In short, these approaches select a single encoder and add new abilities through further pretraining. In contrast, our approach leverages multiple  off the shelf audio encoders that were trained on different domains like general audio(sound), speech etc. and with diverse training mechanisms, including self-supervised learning with training objectives to capture fine-grained information (Huang et al., 2022; Ahmed et al., 2024; Chen et al., 2024; Alex et al., 2025), language aligned contrastive training (Elizalde et al., 2023; Wu et al., 2023; Ghosh et al., 2025b), and transcription based next text token prediction (Radford et al., 2023). Reusing such pretrained encoders avoids redundant pretraining, promotes reuse across communities working on self supervised learning, speech recognition, and CLAP driven multimodal alignment, and improves efficiency in audio LLM development.
When it comes to the integration of audio encoders with the LLM, two architectural paradigms dominate today. The first transforms the outputs of an audio encoder or encoders into the LLM input space (e.g., via an MLP, a QFormer (Li et al., 2023), etc.), then prepend or insert these audio tokens to the text tokens and propagates the entire sequence through all LLM layers as if decoding jointly over audio and text. Please note that the common theme in this family is how audio tokens are passed to the LLM: they are prepended to the text tokens. We refer to this generic scheme Prepend to the LLM’s input token space (PLITS) integration, a term we have introduced to group many state of the art methods in this family of audio LLMs such as  Wu et al. (2025b); Xu et al. (2025b); Chu et al. (2024b); Goel et al. (2025); Chu et al. (2023a); Ghosh et al. (2024); Tang et al. (2024); Gong et al. (2024); Deshmukh et al. (2023). The second paradigm, Flamingo style architectures (Alayrac et al., 2022; Kong et al., 2024), instead insert cross attention and feedforward (FFN) blocks between successive LLM layers; at each insertion, text tokens attend to a set of latent audio tokens, pass through the block FFN, and only then proceed to the next LLM layer. While this design improves attention efficiency relative to PLITS concatenation, the interleaved cross attention plus FFN stacks increase sequential depth and per layer compute, which can slow the forward pass.
In contrast, we introduce LAL, a lightweight integration that injects audio tokens into the LLM’s attention blocks as keys and values only (without forming audio queries) and bypasses the LLM FFNs for audio tokens. This reduces the attention complexity from $\mathcal{O}\!\big((N_{a}{+}N_{t})^{2}\big)$ to $\mathcal{O}\!\big((N_{a}{+}N_{t})N_{t}\big)$, where $N_{a}$ and $N_{t}$ denote the numbers of audio and text tokens, respectively. Since typically $N_{a}\gg N_{t}$, this yields substantial efficiency gains. By avoiding both quadratic attention over audio tokens and their passage through the LLM FFNs, LAL achieves significant reductions in memory usage and computation. Unlike parameter efficient training methods such as LoRA, this is a core architectural modification, so the efficiency benefits are realized not only during training but also at inference time.
Figure 1: Comparison of compute efficiency between LAL (ours) and PLITS, state of the art audio-LLM integration(our baseline). Training was performed with batch size 8 on an NVIDIA A100 using bfloat16, and inference with batch size 12 on an NVIDIA A100 using float16. All benchmarks were executed sequentially on the same node to eliminate load-related discrepancies.
PLITS and Flamingo integration techniques represent complementary strategies for extracting information from audio encoders. LAL provides a compute and memory efficient mechanism by limiting how audio tokens interact with the LLM, while other encoders may benefit from the richer decoding that occurs within the LLM under PLITS style integration. In particular, encoders trained with language contrastive or self supervised objectives such as CLAP and SSLAM are better served by LAL integration, whereas Whisper, which is pretrained with an autoregressive speech that is spoken language transcription and next token prediction objective, gains from the additional decoding capacity of PLITS style integration. Motivated by this observation, we propose a hybrid LAL plus PLITS framework called PAL for building general purpose audio, music, and speech LLMs, enabling encoder aware fusion that balances efficiency with performance. This design achieves strong results while substantially reducing computational and memory requirements compared to using PLITS style integration alone.
To validate these architectural choices, we conduct a systematic empirical study under a standardized training curriculum and dataset setup, ensuring fair comparisons across models. Our experiments explore the trade-off between performance and efficiency, highlighting how encoder-aware fusion facilitates effective information transfer from audio encoders to LLMs with minimal parameter overhead. This analysis provides actionable insights into the design of scalable and efficient audio LLMs that leverage diverse pretrained audio encoders.
Our main contributions are as follows:
1. We introduce LAL, a lightweight integration strategy for audio-LLMs that incorporates audio tokens solely as keys and values in the LLM’s attention sub-modules and skips FFNs, thereby reducing computation and memory cost while retaining performance comparable to PLITS integration.
2. We design an encoder-aware integrated LLM PAL that selectively applies LAL or PLITS integration depending on the audio encoder, enabling general-purpose audio, speech, and music LLMs that balance efficiency with performance.
3. We conduct fair and rigorous architectural comparisons under a standardized training curriculum and dataset setup, providing actionable insights into the efficiency–performance trade-offs of audio-LLM design.
2 Literature review
Audio LLM architectures: When integrating audio encoders with an LLM, two paradigms dominate. In PLITS, encoder features are mapped to the LLM token space with a small projector such as an MLP or a Q Former, the resulting audio tokens are typically prepended to the text tokens, and the joint sequence is processed by all LLM layers (Wu et al., 2025b; Xu et al., 2025b; Chu et al., 2024b; Goel et al., 2025; Chu et al., 2023a; Ghosh et al., 2024; Tang et al., 2024; Gong et al., 2024; Deshmukh et al., 2023). In contrast, the Flamingo style architecture inserts cross attention and feed forward adapters between successive LLM layers so that text tokens attend to latent audio tokens at selected depths (Alayrac et al., 2022; Kong et al., 2024). This makes audio to text interaction explicit and gated, but adds sequential depth, per layer compute, and parameters.
Audio-LLM Datasets: Beyond architecture, recent works curate high-quality instruction tuning datasets, both open source and proprietary (Goel et al., 2025; Ghosh et al., 2024; Chu et al., 2024b; Xu et al., 2025a) and build audio reasoning benchmarks (Sakshi et al., 2024; Deshmukh et al., 2025a; b). Training PLITS or Flamingo style models on these resources improves instruction following and scales reasoning across diverse audio tasks, with most gains attributable to these data innovations.
3 Methodology
This section outlines our approach to integrating audio with language models. We begin by formalizing PLITS, the SOTA audio-LLM integration, as our reference baseline. We then introduce LAL, a lightweight alternative that injects audio through attention only, and we analyze its compute and memory profile. Next, we present the experimental setup and results for classification, captioning, and reasoning, including scaling and a frozen FFN variant. Finally, we connect these findings to PAL, an encoder aware hybrid that selects between PLITS and LAL on a per encoder basis in order to support speech understanding without sacrificing efficiency on general audio.
3.1 Terminologies
We summarize the key terminologies used throughout the paper.
We use PLITS, the SOTA integration as our baseline, LAL as our proposed method, and PAL as a hybrid of the two. Please note that we use  LAL and  PAL to denote the integration approach and the corresponding audio-LLM. We employ SSLAM and CLAP, using an efficient Q-former-based connector that combines information from both inspired by  Tong et al. (2024) without increase in token count, referred to as LFST. When LFST is not used, the audio encoder defaults to SSLAM; when LFST is used, it represents a combination of SSLAM and CLAP. See Appendix E.1 for further details on LFST. Unless otherwise specified, we use Llama 3.2 1B Instruct (Grattafiori et al., 2024) as the base LLM. For evaluating larger models, we report Llama 3.2 3B Instruct  (Grattafiori et al., 2024) , and to assess transfer across model families, we also include Qwen2.5 1.5B Instruct (Team, 2024) .
3.2 Baseline Audio LLM: Prepend to the LLM’s input token space (PLITS)
To establish a fair comparison point for our integration methods, we construct a baseline audio LLM that follows the widely adopted SOTA integration Prepend to the LLM’s input token space (PLITS) paradigm. In this design, the audio encoder outputs are first mapped into the LLM input embedding space using a Q-Former–style connector. The resulting audio tokens are then prepended to the text tokens, and the concatenated sequence is propagated through all LLM layers so that decoding proceeds jointly over audio and text (see Fig. 2(A)). The central characteristic of this paradigm is how audio tokens are provided to the LLM: they are prepended to the text tokens. This integration strategy is used by most audio LLMs, including several state of the art systems Wu et al. (2025b); Xu et al. (2025b); Chu et al. (2024b); Goel et al. (2025); Chu et al. (2023a); Ghosh et al. (2024); Tang et al. (2024); Gong et al. (2024); Deshmukh et al. (2023).
3.3 LAL: Lightweight Audio-LLM Integration
Recent work in mechanistic interpretability suggests that LLMs encode semantics as features that can be selectively activated within hidden states (Elhage et al., 2022; Bricken et al., 2023; Templeton et al., 2024). Building on this view, we hypothesize that effective audio LLM integration requires audio tokens to trigger the activation of sound related conceptual features inside the textual token embeddings. In other words, distinct auditory inputs should induce the corresponding linguistic concepts to become active in the text representation; for example, when the input contains a dog bark, the features associated with the concept dog should light up so the model can ground the auditory signal in language and answer queries such as Which animal sound is present?. This hypothesis guides our architectural design: we seek the simplest pathway that reliably transmits audio cues into the text features that carry concepts.
A standard LLM layer contains an attention submodule followed by a feed-forward(FFN) submodule. Because attention mediates inter token interaction, it is the necessary conduit for audio to influence text, and we posit it is also sufficient for text tokens to gather information from audio. Guided by this principle, we introduce LAL (Lightweight Audio LLM integration). A shared Q Former produces a sequence of audio tokens as in our baseline, and at each layer a MLP projects these tokens into that layer’s input space. Audio information is then injected into the attention block only through Keys and Values while Queries remain text only, so audio modulates the attention context of text tokens without passing through the feed forward network.
Formally, let $H_{l}^{t}\in\mathbb{R}^{N_{t}\times d}$ denote the text hidden states at layer $l$ and $A\in\mathbb{R}^{N_{a}\times d_{a}}$ the Q-Former audio features. A per-layer projector $P_{l}:\mathbb{R}^{d_{a}}\!\to\!\mathbb{R}^{d}$ maps audio to the layer space,
$\hat{A}_{l}=P_{l}(A)\in\mathbb{R}^{N_{a}\times d}$ (1)
and we concatenate text and audio along the token axis
$S_{l}=\big[\,H_{l}^{t}\,;\,\hat{A}_{l}\,\big]\in\mathbb{R}^{(N_{t}+N_{a})\times d}.$ (2)
Queries are formed from text only (see Figure 2(B)), while Keys and Values are computed from the concatenated sequence:
$Q_{l}^{t}=H_{l}^{t}W_{Q,l},\qquad K_{l}=S_{l}W_{K,l},\qquad V_{l}=S_{l}W_{V,l}.$ (3)
The resulting LAL update for text tokens is
$\tilde{H}_{l}^{t}=\mathrm{softmax}\!\left(\frac{Q_{l}^{t}K_{l}^{\top}}{\sqrt{d_{k}}}\right)V_{l}.$ (4)
after which $\tilde{H}_{l}^{t}$ proceeds through the FFN with the usual residual connections. In this way, audio cues shape the attention context seen by text tokens, aligning audio-evoked features with their linguistic counterparts and enabling effective cross–modal information transfer.
Compute and memory efficiency.
LAL improves efficiency over PLITS and Flamingo style along three axes, and the benefits grow with longer audio sequences. We observe up to 64.1% lower memory usage and up to 247.5% higher training throughput (samples/sec). See Figure 1 for detailed training and inference metrics.
A. Attention complexity.
PLITS: full causal attention over $N_{a}+N_{t}$ tokens with cost $\mathcal{O}\!\big((N_{a}{+}N_{t})^{2}\big)$
LAL: only text tokens issue queries; keys and values include audio and text, with cost $\mathcal{O}\!\big((N_{a}{+}N_{t})N_{t}\big)$ eliminating the $N_{a}^{2}$ term and all audio to audio interactions.
B. Feedforward routing.
PLITS: audio tokens pass through attention and the feedforward sublayer in every block, increasing floating point operations and activation memory in proportion to $N_{a}$.
LAL: audio tokens do not enter the feedforward sublayer and only serve as keys and values for text queries, which reduces per layer floating point operations and the activations stored for backpropagation.
Scaling with audio length. Non text modalities in multimodal LLMs often yield far more tokens, and audio is no exception. As $N_{a}$ grows due to longer clips or denser tokenization, PLITS incurs a cost of $(N_{a}+N_{t})^{2}$, so the $N_{a}^{2}$ term dominates. In contrast, LAL scales as $(N_{a}+N_{t})N_{t}$, which is linear in $N_{a}$. Thus the compute and memory gap widens with longer or more finely segmented audio. The feedforward savings in LAL also increase with $N_{a}$ because a larger share of tokens bypass the most expensive part of each block.
Not PEFT or LoRA. (Hu et al., 2022)
LAL is a core architectural change, not a parameter efficient fine tuning(PEFT) method. Techniques such as LoRA modify how weights are adapted during training while keeping the forward compute pattern essentially the same at inference. LAL changes attention and feedforward routing, so its compute and memory savings hold at inference as well as during training.
LAL Integration with Frozen LLM FFN. We also verify that LAL integration remains effective when the LLM’s FFN blocks are frozen, with no significant loss in performance (refer to Appendix E). This finding has important implications for reducing training cost, improving parameter efficiency, and preserving the pretrained knowledge of the LLM while enabling multimodal alignment. For clarity and consistency, however, our main experiments focus on the standard setting with trainable FFN blocks, and discussion of the frozen-FFN variant is limited to Appendix E.
Leveraging parametric versus contextual knowledge: Here we posit how LAL efficiently utilizes two types of knowledge inherent in pre-trained LLMs: (1) parametric knowledge, primarily embedded within the FFN layers as a result of extensive language pre-training, and (2) contextual knowledge, which is dynamically incorporated through attention mechanisms. The empirical success of LAL(refer to Table 1, Table  2) shows that audio input, as contextual information, can effectively induce required concept activations in text token representations via attention-based modulation, without needing direct FFN processing of audio representations. Consequently, audio information indirectly accesses the LLM’s parametric knowledge: the audio context ”piggybacks” on text tokens, as attention mechanisms reconfigure these representations, which then engage relevant concept-related pathways during FFN processing. Such a strategy offers gains in architectural efficiency and provides deeper insights into multimodal information integration.
Figure 2: Illustration of integration techniques: (A) SOTA integration PLITS (prepend to the LLM’s input token space), which prepends audio tokens to text tokens and propagates the full sequence through all LLM layers (our baseline); (B) our proposed lightweight integration LAL, which introduces audio representations only through the attention mechanism (see Equations 2, 3, and 4) while bypassing the feedforward modules; (C) the hybrid PAL, an encoder aware integration that combines LAL and PLITS by selecting the method for each encoder.
Empirical Evaluation of LAL We train and evaluate LAL on general audio tasks, including classification, captioning, and reasoning, across multiple base LLMs following the protocol in Section 4.1. To clearly separate contributions, we present two sets of results. First, in Table 1 (classification and captioning) and Table 2 (reasoning), we report a controlled comparison between LAL and PLITS, showing that LAL achieves comparable or better accuracy while being more efficient in speed and memory. Second, in Table 3 (classification and captioning) and Table 4 (reasoning), we compare LAL with prior works. Note that training data scale and model size vary significantly across prior approaches; our model operates on the lower end of both dimensions. These results should therefore be interpreted as evidence that LAL remains competitive despite using fewer resources.
Table 1: Performance evaluation of the proposed efficient integration method LAL and SOTA integration PLITS across different base LLMs. Evaluation follows the protocol of Gong et al. (2024). AC: Audio caps, CL:Clotho AS2M: AudioSet 2M indicates CIDEr and indicates SPICE. Other metrics: accuracy (ESC-50, VocalSound), Mi-F1 (DCASE), and mAP (FSD, AudioSet). Complete evaluation methodology explained in Section 4.1 and dataset details in Appendix D
LLM PLITS LAL LFST Classification Captioning
Backbone ESC50 DCASE VS FSD AS2M AC CL AC CL
Llama3.2-1B 64.45 37.69 51.57 25.23 9.08 0.59 0.34 16.30 10.96
76.70 40.97 60.87 31.44 11.83 0.66 0.38 16.97 11.87
84.10 45.28 57.59 42.49 14.74 0.70 0.39 17.90 11.82
87.40 46.23 56.03 43.91 14.74 0.72 0.42 18.08 12.58
Llama3.2-3B 70.40 40.62 61.40 28.88 10.84 0.63 0.35 16.81 11.35
82.15 43.21 65.78 34.29 12.91 0.67 0.38 17.80 12.18
89.25 47.21 60.46 43.86 15.03 0.73 0.40 18.61 12.46
Qwen2.5-1.5B 68.00 37.57 56.45 27.87 9.56 0.63 0.38 16.63 11.74
70.85 38.79 59.20 28.53 10.28 0.63 0.38 16.65 11.44
87.80 45.52 56.73 43.26 13.92 0.73 0.41 18.45 12.20
Table 2: GPT-4 evaluation of LAL and PLITS on the CompA-R benchmark (Ghosh et al., 2024). A text only GPT-4 judge scores the model outputs; see Ghosh et al. (2024) for the detailed prompt.
PLITS LAL LFST Helpfulness Clarity Correctness Depth Engagement
3.86 4.74 3.84 2.86 2.99
3.85 4.70 3.82 2.88 3.01
Table 3: Comparison of LAL classification and captioning performance with prior works. Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs, and stronger audio encoders.
Models Classification Captioning
ESC50 DCASE VS FSD AS2M AC CL AC CL
Pengi-124M 91.9 33.8 60.3 46.7 - - - - -
SALMONN-7B 16.4 18.0 16.9 22.1 13.4 - - 8.3 7.6
Audio Flamingo-2-3B 83.9 - -  47.9 - 0.58  0.46 - -
LTU-7B 83.1 45.9 55.6 46.3 18.7 - - 17 11.9
GAMA-7B 82.6 38.4 52.4 47.8 19.2 - - 18.5 13.5
LAL-1B (Ours) 87.40 46.23 56.03 43.91 14.74 0.72 0.42 18.08 12.58
LAL-3B (Ours) 89.25 47.21 60.46 43.86 15.03 0.73 0.40 18.61 12.46
Table 4: LAL performance comparison with prior works for the reasoning (CompA-R) task. All prior works use PLITS integration. Their higher scores mainly stem from larger datasets, bigger LLMs, and stronger audio encoders.
Models Clarity Correctness Engagement Avg
Qwen-Audio-Chat-8B (Chu et al., 2023b) 3.5 3.3 3.6 3.5
LTU-7B (Gong et al., 2024) 3.5 3.2 3.4 3.4
SALMONN-7B (Tang et al., 2024) 2.6 2.4 2.0 2.3
Pengi-124M (Deshmukh et al., 2023) 1.8 1.5 1.3 1.5
LTU w/ CompA-R-7B (Gong et al., 2024) 3.5 3.2 3.4 3.6
GAMA-IT-7B (Ghosh et al., 2024) 4.3 3.9 3.9 4.0
LAL-1B (Ours) 4.70 3.82 3.01 3.80
3.4 PAL: An Encoder Aware Architecture Extending LAL with Speech Understanding
Building on the LAL analysis and results, we ask when PLITS should be preferred over LAL. For the Whisper speech encoder (Radford et al., 2023), our initial experiments on emotion recognition and gender classification indicate that Whisper benefits from decoding inside the LLM; see Table 5. This aligns with classical neuro linguistics where Wernicke’s area is primarily involved in comprehension and has long been associated with processing language in both written and spoken forms, while the angular gyrus supports association across auditory, visual, and other sensory inputs. By analogy, speech features may become most useful when interpreted within a language context, whereas general audio benefits from a modality specific pathway.
Motivated by this, we introduce PAL (Probing the Audio Encoders via LLM), an encoder aware hybrid that chooses the integration per encoder. General audio encoders, SSLAM and CLAP, use LAL integration. The speech encoder, Whisper, uses PLITS integration (refer to Fig. 2 (C)).
Table 5: Integration choices for Whisper evaluated on IEMOCAP (Busso et al., 2008) (emotion recog.) and VoxCeleb2 (Hechmi et al., 2021) (gender cls.) (accuracy, %).
SSLAM+CLAP Integration Whisper Integration IEMOCAP Voxceleb2
PLITS PLITS 65.67 96.69
LAL LAL 66.88 97.19
LAL PLITS 68.81 97.99
Empirical Evaluation of PAL. We train PAL on a unified instruction tuning corpus covering speech, music, and general audio, and evaluate it on classification and reasoning benchmarks. As shown in Table 5 for classification and Tables 6 and 7 for reasoning, PAL is comparable to PLITS in accuracy while retaining efficiency advantages. We also observe that adding a Whisper encoder changes performance in the general audio(sound) and music domains. We hypothesize that this is because Whisper encodes background sounds, as reported by Gong et al. (2023a), which provides some event detection capability.
Our PAL versus PLITS comparison is controlled within our setup, using the same backbone, data, and training hyperparameters; see Appendix C.2 for details. The primary comparison in these tables is therefore between PAL and PLITS, and results from prior work are included only to place PAL in the broader literature. With the exception of Audio Flamingo 2, the other systems are based on PLITS. The higher scores reported by some prior systems over PLITS largely reflect larger training sets, larger LLMs, and stronger audio encoders. This work assesses the integration in isolation, which is why we focus on PAL versus PLITS comparison.
Table 6: Evaluation on MMAU-v05.15.25 (Sakshi et al., 2024) (accuracy, %). Sound (Sn), Music (Mu), Speech (Sp), and overall Average. Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs, and stronger audio encoders. Boldface is used only for fair comparisons.
Sn Mu Sp Total (Avg)
Model mini test mini test mini test mini test
Step-Audio-2-mini-8.3B (Wu et al., 2025a) 79.30 75.57 68.44 66.85 66.18 66.49 72.73 70.23
DeSTA2.5-Audio-8B (Lu et al., 2025) 70.27 66.83 56.29 57.10 71.47 71.94 66.00 65.21
SALMONN-13B (Tang et al., 2024) 41.14 42.10 37.13 37.83 26.43 28.77 34.90 36.23
GAMA-7B (Ghosh et al., 2024) 31.83 30.73 17.71 17.33 12.91 16.97 20.82 21.68
GAMA-IT-7B (Ghosh et al., 2024) 30.93 32.73 26.74 22.37 10.81 11.57 22.83 22.22
LTU-7B (Gong et al., 2024) 20.42 20.67 15.97 15.68 15.92 15.33 17.44 17.23
Qwen2.5-Omni-7B (Xu et al., 2025a) 78.10 76.77 65.90 67.33 70.60 68.90 71.50 71.00
Qwen2-Audio-Instruct-7B (Chu et al., 2024a) 67.27 61.17 56.29 56.29 55.67 55.57 59.90 57.40
M2UGen-7B (Liu et al., 2024b) 43.24 42.44 37.13 38.53 35.37 35.77 37.90 39.76
MusiLingo-7B (Deng et al., 2024) 43.24 41.93 40.12 41.23 31.23 31.73 38.10 38.29
Audio Flamingo-3-8.2B (Goel et al., 2025) 79.58 75.83 73.95 74.47 66.37 66.97 73.30 72.42
Audio Flamingo-2-3B (Ghosh et al., 2025a) 71.47 68.13 70.96 70.20 44.74 44.87 62.40 61.06
Audio Flamingo Chat-1B (Kong et al., 2024) 25.3 23.33 17.66 15.77 6.91 7.67 16.60 15.59
PLITS-1B (Baseline) 71.17 72.20 71.56 69.66 53.45 54.31 65.40 64.61
PAL-1B (Ours) 72.07 70.63 70.66 66.10 53.45 53.28 65.40 63.45
Table 7: Evaluation of PAL on  MMAR (Ma et al., 2025) (accuracy, %). Abbr: Sound (Sn), Music (Mu), Speech (Sp), and Total. Except for Audio Flamingo 2, all other systems use PLITS; their higher scores mainly stem from larger datasets, bigger LLMs, and stronger audio encoders. Boldface is used only for fair comparisons.
Models Sn Mu Sp Mix Mix Mix Mix Total
Sn-Mu Sd-Sp Mu-Sp Sn-Mu-Sp Accuracy
Audio Flamingo-2-3B 24.85 17.48 20.75 18.18 26.61 23.17 8.33 21.90
Audio Flamingo-3-8.2B - - - - - - - 58.5
LTU-7B 19.39 19.90 13.95 18.18 24.77 21.95 16.67 19.20
SALMONN-13B 30.30 31.07 34.69 9.09 34.86 35.37 41.67 33.20
GAMA-7B 29.09 24.27 27.89 27.27 24.77 28.05 20.83 26.50
GAMA-IT-7B 22.42 16.02 12.24 36.36 22.48 14.63 12.50 17.40
Qwen2.5-Omni-7B 58.79 40.78 59.86 54.55 61.93 67.07 58.33 56.70
PLITS-1B (Baseline) 38.79 42.72 40.48 18.18 44.50 39.02 41.67 41.20
PAL-1B(Ours)  40.61 41.75 38.10 36.36 45.87 52.44 41.67 42.20
4 Experimental Setup
4.1 LAL: Experimental Setup
Training Protocol. We train the proposed audio LLM variants on the one of the largest general audio instruction tuning datasets OpenAQA dataset (Gong et al., 2024) and CompA-R Ghosh et al. (2024). Our two-stage pipeline comprises: (i) connector pretraining, where only the connector is trained and all other modules are frozen; and (ii) joint training of the connector and the LLM. The audio encoders remain frozen throughout.
For reasoning and open ended question answering we additionally train on open ended data from OpenAQA Gong et al. (2024) as Stage 3 and on the reasoning dataset CompA R Ghosh et al. (2024) as Stage 4. Additional training details are in Appendix C.1.
Evaluation Protocol. To assess how effectively LAL transfers critical audio-event information from the encoder to the LLM’s latent space, we evaluate on downstream classification, captioning, and reasoning tasks. Following the LTU framework (Gong et al., 2024): (i) for classification, we measure semantic similarity by encoding both model text outputs and target audio labels with gpt-text-embedding-ada; (ii) for captioning, we use standard audio captioning datasets and report CIDEr and SPICE.
For reasoning, we adopt the compA-R-test and the evaluation protocol of (Ghosh et al., 2024): we prompt a text-only GPT-4 judge with the audio-LLM’s output and auxiliary metadata about the audio events, and obtain scores for Helpfulness, Clarity, Correctness, Depth, and Engagement. Additional evaluation details are in Appendix D.1.
4.2 PAL: Experimental Setup
Training Protocol. PAL follows the same two stage procedure as LAL: (i) connector pretraining, where only the connector is trained and all other modules are frozen; and (ii) joint training of the connector and the LLM. The audio encoders remain frozen throughout. For Stage 1, we construct a mixture from the general audio OpenAQA Stage 1 set, augmented with the OpenASQA (Gong et al., 2023b) Stage 1 split for speech understanding. For Stage 2, we use a curated audio, speech, and music reasoning instruction tuning corpus, specifically a 6M subset of AudioSkills (Goel et al., 2025).
Evaluation Protocol. We first target speech understanding with two tasks: speech recognition and speaker gender classification (using gpt-text-embedding-ada as explained in Section 4.1); These are evaluated after Stage 1 to assess how well the newly added Whisper encoder integrates with the LLM. We then assess general audio, music, and speech reasoning on MMAR and MMAU, which report detailed category wise performance.
5 Conclusion
We introduce LAL, which injects audio only through attention keys and values and skips feedforward processing for audio tokens. This reduces attention interactions and activations, yielding up to 64.1% lower memory and up to 247.5% higher training throughput with comparable performance as PLITS, the SOTA baseline integration on classification, captioning, and reasoning. We also propose PAL, an encoder aware hybrid that uses LAL for SSLAM and CLAP and PLITS for Whisper as it benefits from the decoding inside the LLM. LAL is a core architectural change rather than a parameter efficient fine tuning method, so the efficiency gains hold at inference and during training. For future work, we plan to scale to larger backbones, use higher quality instruction data to improve reasoning, and explore streaming and long context audio.
Acknowledgments
This research was supported by the EPSRC and BBC Prosperity Partnership “AI4ME: Future Personalized Object Based Media Experiences Delivered at Scale Anywhere” (EP/V038087/1). For part of the experiments, we used resources provided by the EuroHPC Joint Undertaking, which granted this project access through a EuroHPC Development Access call to the LEONARDO EuroHPC supercomputer hosted by CINECA (Italy) and to the resources of the LEONARDO consortium.
Ethics Statement
All experiments use publicly available datasets. The proposed approach enables beneficial applications, but it could also be misused, for example to monitor individuals without consent. We acknowledge these risks and will release code and models with care, including clear documentation and use guidance to support responsible research.
Reproducibility Statement
Implementation details are provided in Sections 3.3 and 3.4. Training details appear in Appendix C, and the evaluation protocol is described in Appendix D. Code and pretrained models will be made available upon acceptance.
Appendix A Appendix
Appendix B LLM Usage
Large language models were used only as assistive tools for editing and polishing text. We followed the benchmark protocol of  Ghosh et al. (2024) to rate audio LLM outputs; the GPT based evaluation is part of that benchmark. See Section 4.1 for details. LLMs were not used for model design, data selection, experiment setup, implementation, analysis, or generation of results. All technical content was written and verified by the authors.
Appendix C Training Details
C.1 LAL Training Details
We use OpenAQA (Gong et al., 2024) two stage training setup for LAL to report the results in Table 1. We also train on broader open ended data from OpenAQA (Gong et al., 2024) and on the reasoning dataset CompA R (Ghosh et al., 2024), with evaluations shown in Table 2. Additional training hyperparameters appear in Table 8.
Table 8: Hyper-parameters used for the three stage training of LAL and PLITS (Llama3.2 1B)
Training Configuration Stage 1 Stage 2 Stage 3 $|$ Stage 4
(Connector Pre training) (LLM Fine tuning) (LLM Fine tuning)
Optimizer AdamW (Loshchilov & Hutter, 2017)
Learning Rate Schedule Cosine (Loshchilov & Hutter, 2016)
Peak Learning Rate 0.001 0.0001 0.0001
Epochs 1 1 1
Warm up Ratio (steps) 0.05 0.03 0.03
Dataset Size 1.2 M 1.9 M 5.6 M $|$ 200 K
Batch Size 3232 1212 1212
Gradient Accumulation Steps 4
GPUs $2\times$ Nvidia A100 (80GB)
RAM 150 GB
Loss Next token loss on text part
C.2 PAL Training Details
PAL uses a two stage training protocol(Table  9). In Stage 1, we start from the Stage 1 dataset used for LAL and augment it with additional speech focused data from OpenASQA (Gong et al., 2023b). In Stage 2, we fine tune on a curated audio, speech, and music reasoning instruction corpus, AudioSkills (Goel et al., 2025). We use a 6M example subset of AudioSkills (from the original 10M) due to the unavailability of original audio files for some source datasets.
Table 9: Hyperparameters used for the two stage training of PAL and PLITS (Llama3.2 1B)
Training Configuration Stage 1 Stage 2
(Connector Pre training) (LLM Fine tuning)
Optimizer AdamW (Loshchilov & Hutter, 2017)
Learning Rate Schedule Cosine (Loshchilov & Hutter, 2016)
Peak Learning Rate 0.001 0.0001
Epochs 1 1
Warm up Ratio (steps) 0.05 0.03
Dataset Size 1.7 M 6.4 M
Batch Size 1616 44
Gradient Accumulation Steps 22 3232
GPUs $4\times$ Nvidia A100 (64GB)
RAM 250 GB
Loss Next token loss on text part
Appendix D Evaluation Details
D.1 LAL Evaluation Details
We follow the evaluation protocol of Gong et al. (2024) for classification and captioning, and use the CompA R test set of Ghosh et al. (2024) for reasoning. Below we summarize the datasets included in the Gong et al. (2024) protocol.
VocalSound (Gong et al., 2022b): The VocalSound dataset consists of 21,024 crowd-sourced recordings of 6 different classes of vocal expressions collected from 3,365 unique subjects. We evaluated our model on the VocalSound evaluation set which contains 3,594 audio clips, and report top-1 accuracy scores across the 6 classes for single-class classification performance. It is important to note that VocalSound was excluded from our training data; therefore, our evaluation on VocalSound is considered zero-shot.
ESC-50 (Piczak, 2015): The ESC-50 dataset comprises 2,000 five-second environmental audio clips categorized into 50 different classes. Following Gong et al. (2024), we evaluate our model on all 2,000 audio samples and report the top-1 accuracy score for single-class classification performance. It is important to note that while ESC-50 is originally sampled from the Freesound dataset (which is included in our training data), ESC-50 itself was excluded from training. Therefore, our evaluation on this dataset is considered a weak zero-shot evaluation.
DCASE2017 task 4 (DCASE) (Mesaros et al., 2019): DCASE 2017 Task 4 contains 17 sound events distributed across two categories: ”Warning” and ”Vehicle”. The evaluation set consists of 1,350 audio clips. We evaluated our model on this dataset and report micro F1-score(MiF1) for single-class classification performance. It is important to note that DCASE 2017 task 4 is originally sampled from AudioSet, which is included in our training data. However, DCASE 2017 task 4 itself is excluded from training, making our evaluation on this dataset a weak zero-shot evaluation.
FSD50K (FSD) (Fonseca et al., 2021): The FSD50K evaluation set contains 10,231 audio clips. We evaluated our model on this evaluation set and report the mAP score for multi-label classification performance. Since the training and validation sets of FSD50K are included in our training data, this evaluation is considered an in-domain evaluation.
AudioSet (Gemmeke et al., 2017): We evaluated our model on this evaluation set and report the mAP score for multi-label classification performance. The training set of AudioSet is included in our training data, making this evaluation an in-domain evaluation.
AudioCaps (Kim et al., 2019): The AudioCaps evaluation set contains 901 audio clips, each paired with 5 audio captions, resulting in a total of 4,505 audio-caption pairs. We evaluated our model on this evaluation set and report the captioning scores using CIDER and SPICE metrics. The training and validation sets of AudioCaps are included in our training data, making this evaluation an in-domain evaluation.
Clotho V2 (Drossos et al., 2020): The Clotho V2 evaluation set contains 1,045 audio clips, each paired with 5 audio captions, resulting in a total of 5,225 audio-caption pairs. We evaluated our model on this evaluation set and report the captioning scores using CIDER and SPICE metrics. The development and validation sets of Clotho V2 are included in our training data, making this evaluation an in-domain evaluation.
D.2 PAL Evaluation Details
For speech classification (emotion recognition and gender classification), we follow the protocol of Gong et al. (2023b). For combined sound, speech, and music reasoning, we evaluate on the standard benchmark datasets MMAU (Sakshi et al., 2024) and MMAR (Ma et al., 2025).
Appendix E LAL Integration with Frozen LLM FFN
Standard audio-LLM training typically requires full fine tuning of the LLM. However, since LAL integrates audio information solely through the attention mechanism, we investigate whether LAL remains effective when the LLM feedforward (FFN) blocks, which are widely believed to encode much of the model’s factual and linguistic knowledge, are frozen and only the attention layers are updated. In Stage 2 of our training pipeline, we therefore construct a variant with the LLM FFN frozen. As shown in Table 10, performance is largely maintained under this setting. This result suggests that LAL can successfully integrate audio information through attention without modifying the knowledge stored in the FFN modules. Such a property has important implications for reducing training cost, improving parameter efficiency, and preserving the pretrained knowledge of the LLM while enabling multimodal alignment.
Table 10: Performance evaluation of the LAL Integration with frozen FFN. Evaluation follows the protocol of Gong et al. (2024). AC: Audio caps, CL:Clotho AS2M: AudioSet 2M indicates CIDEr and indicates SPICE. Metrics: accuracy (ESC-50, VocalSound), Mi-F1 (DCASE), and mAP (FSD, AudioSet). Complete evaluation methodology explained in Section 4.1 and dataset details in Appendix D
LLM FFN PLITS LAL LFST Classification Captioning
Backbone Frozen ESC50 DCASE VS FSD AS2M AC CL AC CL
Llama3.2-1B 64.45 37.69 51.57 25.23 9.08 0.59 0.34 16.30 10.96
 76.70  40.97  60.87  31.44  11.83  0.66 0.38  16.97  11.87
71.80 33.99 55.28 29.38 10.48 0.63  0.40 16.11 11.75
E.1 LFST Connector: Language aligned and Fine grained Spatiotemporal Connector
We adopt the connector proposed in Cambrian (Tong et al., 2024) and apply it in our audio setting to fuse a language aligned encoder such as CLAP with a self supervised encoder such as SSLAM. The connector produces a compact set of latent tokens that combine semantic cues from CLAP with fine grained spatiotemporal detail from SSLAM, while keeping sequence length fixed and avoiding the overhead of naive concatenation.
Formalization.
Let the encoder outputs be
$H_{\text{sslam}},\;H_{\text{clap}}\in\mathbb{R}^{F\times T\times d},\quad z\in\mathbb{R}^{d},$
where $F$ is frequency, $T$ is time, and $d$ is the feature dimension. Following Tong et al. (2024), a single latent token $z$ is broadcast to each spatiotemporal location, yielding $z_{f,t}$ for every $(f,t)$. Inside the connector, which consists of 3 cross attention layers, each $z_{f,t}$ is updated through cross attention with the corresponding local regions of $H_{\text{sslam}}$ and $H_{\text{clap}}$. To preserve temporal structure when flattening across $(F,T)$, we insert a newline token along the frequency axis so that each new time step begins with this marker before its spectral tokens (see Figure 3).
Figure 3: Overview of LFST using the Cambrian connector (Tong et al., 2024). A single latent token is broadcast to every time–frequency location and then updated inside the connector by cross attention with local SSLAM and CLAP features, fusing fine grained spatiotemporal detail with language aligned semantics. The red tokens illustrate the latent query and the local encoder keys and values it attends to. A newline token is inserted at each new time step so the flattened sequence preserves the original spatiotemporal layout while keeping the output length fixed.